OpenStudy (anonymous):

A funnel is shaped approximately like a cone. The diameter of the base is 6 inches. The height is 10 inches. What is the approximate volume of the funnel? Use 3.14 for pi, and round to the nearest hundredth. a. 94.2 in.3 c. 226.08 in.3 b. 18.84 in.3 d. 28.26 in.3

OpenStudy (anonymous):

The volume of a cone is: Vcone = 1/3 × pi × r^2 × h = 1/3 * 3.14 * (6/2)^2 * 10 = 94.2 in.3. The approximate volume of the funnel is 94.2 cubic inches.